OpenStudy (anonymous):

An ocicat eats \frac 35 of a pound of food daily. How many ocicats can a 19\frac 12-pound bag of food feed for one week

OpenStudy (help!!!!):

I don't understand the /frac 35?

OpenStudy (anonymous):

/frac 35?=3/5 19\frac 12 = 19 1/2

OpenStudy (help!!!!):

yeah so just \[\frac{ 3 }{ 5 }\div \frac{ 39 }{ 2 }\]

OpenStudy (help!!!!):

because 3/5 is 1 day and if we need to find out how many it consumes in a week or 7 days

OpenStudy (help!!!!):

do you understand?

OpenStudy (help!!!!):

3/5 is one cat so divide by the amount of food or 19 1/2 to find out how many cats it can feed @jenise